In a circle whose radius is 6 cm, there is a central angle whose measure is 60 degrees.How long is the chord joining the endpoints of the arc cut off by the angle?
Answer by jim_thompson5910(35256)

This is an isosceles triangle since the two radii (that are part of this triangle) are congruent legs.

So the base angles are both congruent to x. We have the 3 angles: 60, x, x

They must add to 180, so...

60 + x + x = 180

60 + 2x = 180

2x = 180 - 60

2x = 120

x = 120/2

x = 60

This means that all three angles of this triangle are 60 degrees. So we actually have an equilateral triangle with all sides equal to 6 cm, which is the final answer.